# Birgir Hrafnkelsson Morris Riedel Egill Skúlason

### v1

@@ 8.4 Farthest point sampling
Figure 33 demonstrates the practical advantage of this approach in the singlet_016 system, where FPS and adaptive variance control enable robust and efficient saddle searches, in contrast to the instabilities observed in standard GPDimer runs, while also mitigiating the effect of the size of the hyperparameter matrix optimization, which is the primary walltime bottleneck.
Figure 33: Performance trace for the singlet_016 system (Figure fig:equiv:optgd), illustrating the comparative behavior of GPDimer and OTGPD during saddle search optimization. (A) The per-iteration electronic structure function counts and wall time show that OTGPD (skyblue) consistently achieves lower and more stable computational cost per iteration compacoral to GPDimer (coral), which exhibits pronounced spikes and variability. (B) Convergence profiles of the maximum force component (log scale) demonstrate smoother and more rapid relaxation for OTGPD, while GPDimer progress stalls intermittently, reflecting underlying model instability. (C) Evolution of key hyperparameters over the course of the optimization, with the GP signal variance (magenta, Var) and interatomic distances (C-C, H-C, H-O, C-O, H-H, O-O) tracked for both methods. GPDimer displays episodes of pathological variance growth, coinciding with force and runtime spikes, whereas OTGPD maintains stable and physically reasonable hyperparameter values throughout.
@@ 8.5 Variance control and Hyperparameter stability
Figure 33 also shows a sudden jump in the hyperparameters, in particular, the variance. The effect of this hyperparameter is to basically make the model pathlogoically sensitive to the data points, which in turn will sample configurations which crash NWChem. To counteract this instability and ensure robust performance, we implement two complementary mechanisms discussed in chapter 7: a direct control on the signal variance via an adaptive barrier, and a general heuristic for monitoring the stability of all hyperparameters. The adaptive behavior eliminates pathological variance growth while preserving the surrogate’s ability to capture the true curvature of the Potential Energy Surface. Empirically, the combination of the adaptive barrier on $\sigma_{f}^{2}$ and the general HOD mechanism provides robust control. The barrier acts as a targeted preventative measure against a known failure mode, while the HOD acts as a general safety net for the entire optimization process. Together, they reduce the number of failed saddle searches from roughly twelve percent to two percent, contributing significantly to the overall efficiency and reliability of the OTGPD method.
@@ 8.6.2 Linear bending angles and Sella
Figure 35: Endpoints for saddle point search trajectories of the singlet_016 system starting from an initial configuration (A). The standard Dimer method (B) and the proposed OTGPD method (E) identifies the nearest transition state structure. The previous GPDimer method (D) and Sella (C) is guided towards a much more fractured state.
@@
A two-dimensional landscape projection (Figure 37) maps pathways from the initial, high-energy configuration to the deep intermediate minimum and transition states. The white star marks the OTGPD saddle, which is clearly proximal to the initial point, while the surface shows the trajectory of paths explored during the optimization.
Figure 37: A 2D landscape projection visualizing the potential energy surface of the n-propyl acetate system. This surface, described Sec. 4.2 depicts the energy landscape as a function of observed paths during the optimization. The landscape clearly several states. The proximal transition state (white star), corresponding to C-O cleavage, is the saddle point located by OTGPD and Dimer. The more distant saddle is the endpoint, corresponding to the 1,5-HAT, located by Sella. This visualization strongly suggests that Sella’s trajectory overshoots the first, more proximal saddle. We indicate the unconverged climbing image to note that the dimer saddle cannot be found on the path, though the Sella trajectory passes near the dimer saddle configuration as reported earlier [36].
@@ 9 Summary
This thesis presents a multi-faceted investigation into the role of software representation as a primary driver of progress in computational science. The key contributions are summarized as follows:
First, a high-order finite element solver, featom, was developed for relativistic atomic structure calculations. The success of this tool demonstrated that physical fidelity is achieved through a cascade of representations at the mathematical, numerical, and software levels.
Second, a comprehensive modernization of the EON simulation package was undertaken. This foundational software engineering effort transformed the legacy code into a robust, modern platform. This architectural overhaul was a direct prerequisite for the development of novel scientific methods, including a high-performance client-server architecture for potential evaluation and a new hybrid Climbing Image Nudged Elastic Band (CI-NEB-MMF) algorithm.
Third, a rigorous statistical methodology was applied to the benchmarking of transition-state search algorithms. Using a Bayesian generalized linear mixed-effects model, the performance of optimizers for the dimer method rotation phase was analyzed. This provided statistically robust evidence that the Conjugate Gradient (CG) algorithm is both more efficient and significantly more reliable than the commonly used L-BFGS optimizer for this task.
Finally, the Optimal Transport Gaussian Process (OT-GP) framework was developed to solve the critical wall-time bottleneck in GP-accelerated saddle searches. By introducing a chemically-aware farthest-point sampling strategy for hyperparameter optimization and a suite of novel stability controls, the OT-GP method achieves a 96% success rate and halves the wall-time-to-solution compared to its predecessors. This elevates GP-based acceleration from a promising but fragile technique into a robust and efficient tool for exploring complex chemical landscapes.
@@ 10 Conclusions
As summarized in the preceding chapter, the work of this thesis has produced a suite of novel methods and robust software tools, transforming raw computation into scientific insight. Yet, as the doctrine of Alan Perlis suggests, each solved problem reveals the next. This chapter moves beyond a summary of results to synthesize their broader implications, confront the limitations and open questions that remain, and outline a path forward for the future of computational science.
@@
The development of powerful, specialized tools requires a significant investment of expert effort, yet their long-term maintenance and the management of their inherent technical debt represent a systemic challenge in academic research. The future of the field depends not only on the creation of novel algorithms but on the development of a science of scientific software itself by treating the software itself as a first-class research object, we can ensure that the insights gained from one generation of scientific inquiry provide a robust and lasting foundation for the next.
The central conceptual point in terms of a Gaussian Process stems from the definition, as a MVN over function values, with MLL to select the MVN that best explains the observed data under the Gaussian Process prior, we cannot be guaranteed to reproduce the true Potential Energy Surface $V(x)$ generated by electronic structure calculations. The many body equation [115],
$V(x)=\sum_{i}V_{1}(i)+\sum_{i<j}V_{2}(i,j)+\sum_{i<j<k}V_{3}(i,j,k)+\dots+V_{N}(1,\dots,N),$ (99)
is not generally Gaussian-distributed across geometries, nor is it characterized by the two-point covariances implied by a stationary kernel. While conceptually simple, the note appears easily forgotten, and bears repeating. With sparse, correlated data along a saddle-search path, maximizing MLL admits degenerate directions which is most visibly evident in the signal variance that produce permissive, high-variance posteriors with flattened means.
Across the breadth of this thesis, an attempt has been made to bring to the fore the requirement that the only real way forward towards efficient representations is through the effort expended in “speaking binary.” Formula translation in the exascale era is no longer feasible, as changes in underlying hardware bring additional constraints on algorithms. Despite this, the free lunch is not yet over, the OT-GP is essentially almost an exercise in profiling, that the hyperparameters estimates are the slowest step naturally arises from a study of the traces of accelerated saddle searches. The OT-GP framework, while successful, still shares the same worst-case time complexity as its predecessors; pruning over dynamic data remains an unsolved problem. Technical mastery over the average-time complexity must be accompanied by a proliferation in terms of ease of use, a direction covered in the architectural vision of a “BLAS for chemical kinetics.”
This vision provides the path to scaling the powerful new methods developed herein. The OT-GP framework now provides a foundational approach for the active learning of high-energy transition state geometries, essential for training next-generation reactive machine-learned potentials, with applications in photochemistry and adaptive kinetic Monte Carlo. Future state-of-the-art reductions will likely involve symmetry adaption or reinforcement learning, built upon decoupled, polyglot libraries communicating through zero-copy interfaces.

### v2

@@ 8.4 Farthest point sampling
Figure 35 demonstrates the practical advantage of this approach in the singlet_016 system, where FPS and adaptive variance control enable robust and efficient saddle searches, in contrast to the instabilities observed in standard GPDimer runs, while also mitigating the effect of the size of the hyperparameter matrix optimization, which is the primary walltime bottleneck.
Figure 35: Performance trace for the singlet_016 system (Figure fig:equiv:optgd), illustrating the comparative behavior of GPDimer and OTGPD during saddle search optimization. (A) The per-iteration electronic structure function counts and wall time show that OTGPD (skyblue) consistently achieves lower and more stable computational cost per iteration compared to GPDimer (coral), which exhibits pronounced spikes and variability. (B) Convergence profiles of the maximum force component (log scale) demonstrate smoother and more rapid relaxation for OTGPD, while GPDimer progress stalls intermittently, reflecting underlying model instability. (C) Evolution of key hyperparameters over the course of the optimization, with the GP signal variance (magenta, Var) and interatomic distances (C-C, H-C, H-O, C-O, H-H, O-O) tracked for both methods. GPDimer displays episodes of pathological variance growth, coinciding with force and runtime spikes, whereas OTGPD maintains stable and physically reasonable hyperparameter values throughout.
@@ 8.5 Variance control and Hyperparameter stability
Figure 35 also shows a sudden jump in the hyperparameters, in particular, the variance. The effect of this hyperparameter is to basically make the model pathologically sensitive to the data points, which in turn will sample configurations which crash NWChem. To counteract this instability and ensure robust performance, we implement two complementary mechanisms discussed in chapter 7: a direct control on the signal variance via an adaptive barrier, and a general heuristic for monitoring the stability of all hyperparameters. The adaptive behavior eliminates pathological variance growth while preserving the surrogate’s ability to capture the true curvature of the Potential Energy Surface. Empirically, the combination of the adaptive barrier on $\sigma_{f}^{2}$ and the general HOD mechanism provides robust control. The barrier acts as a targeted preventative measure against a known failure mode, while the HOD acts as a general safety net for the entire optimization process. Together, they reduce the number of failed saddle searches from roughly twelve percent to two percent, contributing significantly to the overall efficiency and reliability of the OTGPD method.
@@ 8.6.1 Reliability
8.6.2 Identifying failure modes
The OTGPD framework successfully eliminates the signal variance instability that plagued earlier iterations. However, our benchmarking revealed four specific systems—D016, D084, D100, and S242—where the OTGPD failed to converge, despite success in the GPDimer baseline. A forensic analysis of these trajectories indicates that these failures stem not from an intrinsic algorithmic instability in the OTGPD, but rather from artifacts introduced by the dimer initialization routine.
For the OTGPD benchmark, we generated initial dimer configurations by displacing atoms along the softest mode identified by an inexpensive semi-empirical (GFN-xTB) calculation. In these four specific cases, this procedure produced pathological starting geometries characterized by excessively high forces or unphysical atomic overlaps. Figure 37 contrasts these starting points. For example, system D016 initialized with unphysically close carbon atoms, while S242 erroneously replaced a hydrogen on the middle carbon with a methyl end-group.
When a GP-accelerated method receives such a “poisoned” baseline—a high-energy, high-force configuration—the initial surrogate model incorporates this unphysical data. The Gaussian Process effectively learns that these extreme forces are characteristic of the landscape, leading to a warped posterior that guides the search into unstable regions. The GPDimer benchmark avoided this specific pathology by utilizing a different initialization strategy based on displacing the least coordinated atom.
Figure 37: Initializations for the GPDimer (top) and OTGPD (bottom). The xTB initialization procedure in these four cases results in unphysical initial geometries, leading to failures in the optimization routine. Specifically, D016 exhibits unphysically close carbon atoms, D084 possesses a shortened carbon-oxygen bond, D100 shows near-overlapping carbon atoms, and S242 features a misplaced methyl group.
To verify this diagnosis, we re-executed the OTGPD searches for these four systems using the valid, non-pathological initial configurations from the GPDimer runs. Table 6 presents the results. In every case, the OTGPD successfully converged. Furthermore, it retained its performance advantage: for the challenging D084 system, OTGPD required only 3.0 minutes (181 s) compared to 6.95 minutes (417 s) for GPDimer and 4.25 minutes (255 s) for the standard Dimer. This confirms that the observed “failures” were initialization artifacts rather than deficiencies in the optimal transport framework.
Table 6: Performance metrics for re-run systems (D016, D084, D100, S242) using identical initial configurations. All methods converge successfully. Times reported in minutes. The OTGPD maintains superior or competitive efficiency in both electronic structure calls and total wall time.
System Method HF Calls Time (min)
D016 Dimer 106 7.4
GPDimer 20 3.7
OTGPD 20 3.3
D084 Dimer 2666 255.4
GPDimer 75 417
OTGPD 65 181.4
D100 Dimer 214 24.3
GPDimer 28 16.2
OTGPD 28 18.52
S242 Dimer 249 14.6
GPDimer 25 2.3
OTGPD 33 1.85
@@ 8.6.3 Linear bending angles and Sella
Figure 38: Endpoints for saddle point search trajectories of the singlet_016 system starting from an initial configuration (A). The standard Dimer method (B) and the proposed OTGPD method (E) identifies the nearest transition state structure. The previous GPDimer method (D) and Sella (C) are guided towards a much more fractured state.
@@
A two-dimensional landscape projection [122] (Figure 40) maps pathways from the initial, high-energy configuration to the deep intermediate minimum and transition states. The white star marks the converged OTGPD saddle, which is clearly proximal to the initial point, while the surface shows the trajectory of paths explored during the optimization.
Figure 40: A 2D landscape projection visualizing the potential energy surface of the n-propyl acetate system. This surface, described in Sec. 4.2, depicts the energy landscape as a function of observed paths during the optimization. The landscape clearly reveals several states. The proximal transition state (white star), corresponding to C-O cleavage, is the converged saddle point located by OTGPD and Dimer. The more distant saddle is the endpoint corresponding to the 1,5-HAT, located by Sella. This visualization strongly suggests that Sella’s trajectory overshoots the first, more proximal saddle. We highlight the converged dimer saddle to illustrate that while the Sella trajectory passes near this configuration, it fails to localize the proximal transition state, instead proceeding to the distal 1,5-HAT saddle. The Sella trajectory passes near the dimer saddle configuration as reported earlier [50].
@@ 8.6.4 Performance
8.7 Conclusions
This chapter transformed the Gaussian Process from a volatile statistical interpolator into a robust engine for chemical discovery. We demonstrated that naive application of machine learning to Cartesian landscapes frequently yields pathological failures, such as the “cold fusion” artifacts observed in the standard GPDimer. These failures stem not from a lack of data, but from a lack of physical constraint.
To remedy this, we replaced the arbitrary Euclidean metric with the permutation-invariant Intensive Earth Mover’s Distance (EMD). This metric equips the surrogate with a chemically meaningful sense of locality. It underpins the adaptive trust radius, a mechanism that strictly enforces the validity of the local approximation and prevents the optimizer from wandering into unphysical regions of the potential energy surface. Furthermore, by linking data selection to numerical stability via Farthest Point Sampling, we ensured that the covariance matrices remain well-conditioned, satisfying Gershgorin bounds and enabling reliable inversion even in high-throughput regimes.
Empirical validation confirms the efficacy of this approach. The OTGPD lies on the Pareto-optimal frontier for the majority of test systems, offering the most favorable trade-off between wall-clock time and electronic structure calls. Crucially, the comparison with Sella reveals a distinct topological advantage: while internal-coordinate methods often overshoot to find thermodynamically stable but geometrically distant states, OTGPD reliably converges to the proximal saddle point. This capability proves essential for kinetic Monte Carlo simulations, where discovering the immediate connectivity of the reaction network takes precedence over finding the global minimum.
Ultimately, the OTGPD represents the culmination of the efficiency strategies developed throughout this thesis. It integrates the robust saddle-finding logic of the Dimer method, the statistical power of Gaussian Processes, and the rigorous stability controls derived from optimal transport theory into a single, cohesive framework for autonomous reaction discovery.
@@ 9 Summary
This dissertation presented a multi-faceted investigation into the role of computational representation as a primary driver of progress in chemical physics. We structured this inquiry around specific research objectives targeting relativistic electronic structure, software architecture, statistical benchmarking, and algorithmic efficiency. Here, we summarize the contributions against those initial goals.
We began by questioning whether a squared Hamiltonian formulation within a finite element framework could provide stability for all-electron relativistic calculations. Chapter 3 validated this hypothesis. The featom solver, built upon a high-order FEM discretization of the squared Dirac operator, successfully eliminated spectral pollution. It achieves sub-second wall times for all-electron DFT calculations of heavy elements like Uranium, outperforming state-of-the-art shooting methods in the relativistic regime while avoiding the complexity of kinetic balance constraints.
Subsequently, we investigated whether modernizing legacy software architectures could unlock novel scientific algorithms. Chapter 4 demonstrated this through the complete refactoring of the EON suite. By decoupling the potential energy surface via a client-server architecture, we enabled the creation of the hybrid MMF-NEB method. This algorithm reduces the computational cost of double-ended saddle searches by approximately 60% compared to standard approaches, a gain achievable only through the flexibility of the redesigned state management system. In Chapter 5, a C++ rewrite enabled the Cartesian coordinate based GPDimer to compete with state of the art internal coordinate methods.
To ensure these gains were robust, we sought to move beyond average-cost benchmarks and rigorously quantify algorithmic reliability. Chapter 6 applied Bayesian generalized linear mixed models to the Dimer method. This analysis overturned the conventional wisdom regarding rotational minimization, providing robust statistical evidence that the Conjugate Gradient optimizer offers superior reliability compared to L-BFGS. Furthermore, we quantified the high cost of quaternion-based rotation removal within EON, demonstrating that it incurs a significant computational penalty without a credible improvement in convergence success.
Finally, we addressed the cubic scaling and instability bottlenecks of Gaussian Process acceleration. Chapter 8 introduced the OTGPD framework. By implementing the Intensive Earth Mover’s Distance for physically motivated trust regions and Farthest Point Sampling for numerical conditioning, we eliminated the “cold fusion” pathologies observed in previous methods. The resulting algorithm places on the Pareto frontier for 190 out of 238 test systems, effectively halving the time-to-solution compared to the standard GPDimer and demonstrating superior performance to internal-coordinate methods for proximal saddle searches.
@@ 10 Conclusions
As summarized in the preceding chapter, this thesis produced a suite of novel methods and robust software tools, transforming raw computation into scientific insight. Yet, as the doctrine of Alan Perlis suggests, every solution unmasks a new layer of problems. This chapter moves beyond a summary of specific results to synthesize their broader implications, confront the limitations that remain, and outline a path forward for the future of computational science.
10.1 The Science of Scientific Software
@@
The development of powerful, specialized tools demands a significant investment of expert effort, yet their long-term maintenance and the management of their inherent technical debt represent a systemic challenge in academic research. The future of the field depends not only on the creation of novel algorithms but on the elevation of software construction to the status of a science itself. By treating the software as a first-class research object, we ensure that the insights gained from one generation of scientific inquiry provide a robust foundation for the next.
Across the breadth of this thesis, we emphasized that efficient representations emerge only through the effort expended in “speaking binary.” The era of “formula translation”—blindly transcribing mathematical notation into code—has ended. The constraints of modern hardware, particularly in the exascale era, demand a deeper engagement with memory hierarchies, vectorization, and concurrency. The success of the OTGPD serves as a testament to this philosophy; it functions essentially as an exercise in rigorous profiling. We identified that hyperparameter estimation constituted the wall-time bottleneck and addressed it not through clearer mathematics, but through algorithmic restructuring and data pruning. Thus, technical mastery over average-time complexity must accompany theoretical innovation.
10.2 The Ontological Gap: Statistics vs. Physics
A central tension persists throughout this work: the disconnect between the statistical surrogate and the physical reality it models. A Gaussian Process defines a MVN over function values. We select the specific distribution that best explains the observed data by maximizing the MLL. However, this statistical procedure implies no guarantee that the model reproduces the true Potential Energy Surface generated by electronic structure theory.
The true potential energy surface arises from the many-body expansion [159]:
$V(x)=\sum_{i}V_{1}(i)+\sum_{i<j}V_{2}(i,j)+\sum_{i<j<k}V_{3}(i,j,k)+\dots+V_{N}(1,\dots,N)$ (100)
This expansion possesses specific decay properties and symmetries not inherently captured by the two-point covariances of a stationary Gaussian kernel. The GP “anthropomorphizes” the data; we often ascribe unjustified physical meaning to hyperparameters like the signal variance or length scale. In reality, these parameters merely maximize the likelihood of a specific dataset. When that dataset consists of sparse, correlated points along a saddle-search path—atypical of the global surface—the MLL surface often becomes shallow or degenerate. This manifests visibly in the “cold fusion” pathologies observed in the standard GPDimer, where the signal variance explodes to accommodate high-force gradients, flattening the mean prediction and directing the search into unphysical regions.
The OTGPD mitigates this not by forcing the GP to understand physics, but by imposing physical constraints (trust regions, stability barriers) upon the statistical engine. We acknowledge that the model operates as a mathematical interpolator, not a physical simulator, and constrain it accordingly.
10.3 Future Outlook: A BLAS for Chemical Kinetics
The OT-GP framework, while successful, shares the worst-case cubic time complexity of its predecessors. Pruning over dynamic data remains an open challenge, particularly for systems exceeding the size regimes explored here. To address this, the field requires a paradigm shift in software architecture.
We propose the vision of a “BLAS for chemical kinetics.” Just as the Basic Linear Algebra Subprograms (BLAS) provide a standard, highly optimized foundation for linear algebra, we require a standardized, decoupled layer for kinetic primitives—saddle searches, rate evaluations, and traversing potential landscapes. Such a framework would rely on polyglot libraries communicating through zero-copy interfaces, allowing researchers to compose high-level workflows (like AKMC or MD) from highly optimized, hardware-aware components.
This architectural evolution paves the path for scaling the methods developed herein. The OTGPD provides the foundational approach for the active learning of high-energy transition state geometries. These datasets act as the critical feedstock for training next-generation reactive machine-learned potentials. By integrating these surrogates with symmetry-adapted representations or reinforcement learning agents, we move closer to the ultimate goal: the autonomous, reliable exploration of the vastness of chemical space.